OpenStudy (anonymous):

Identify the values restricted from the domain of the expression. x+6 / x^2 + 7x-8

OpenStudy (cwrw238):

Hint the denominator cannot have a vule of zero

OpenStudy (anonymous):

Can one of the values be 1..?

OpenStudy (cwrw238):

yes that would make the bottom part zero if you factor the denominato you'll find the other value

OpenStudy (anonymous):

I got 8..but I think it might be -8..

OpenStudy (anonymous):

Would that be right?

OpenStudy (cwrw238):

x^2 + 7x - 8 = (x - 8)(x + 1) so if x is either 8 or -1 the denominator will be 0

OpenStudy (cwrw238):

Its -1 not 1

OpenStudy (anonymous):

wouldn't it be 1 and -8?

OpenStudy (anonymous):

as the restricted values?

OpenStudy (cwrw238):

oooh - whats wrong with me today forget that last post and start again x^2 + 7x - 8 = (x + 8)(x - 1) so x = -8 or x = 1 will make the denominator zero the restricted values are x = 1 and -8 right second time lol!

OpenStudy (cwrw238):

restricted values from the domain are 1 and -8.

OpenStudy (cwrw238):

yes - you were right

OpenStudy (anonymous):

Thank you! :D

OpenStudy (cwrw238):

yw